40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15662     Entry time: Fri Nov 6 14:08:44 2020
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC re-locked 

The PMC servo railed and so I re-locked it at ~half range. I've been noticing that the diurnal drift of the PZT control voltage has been larger than usual - not sure if it's entirely correlated with temperature on the PSL table. Anyway the cavity is locked again so all is good.

ELOG V3.1.3-